# A Survey on Speech Large Language Models for Understanding

### v5

@@ VII-A2 Inaccurate Evaluation
In summary, our dual-track evaluation framework accommodates structured non-structured and weakly-structured tasks, fostering comprehensive and principled analysis of Speech LLMs across diverse speech understanding scenarios. To further ground this framework in practice, the following sections present a detailed analysis of representative tasks from each category—namely, ASR as a canonical structured task, ST as a prototypical weakly-structured task, and multitask capabilities as a hallmark of unstructured understanding.
@@ VII-B1 Analysis Setup
RPS=Model ScoreSOTA $\text{RPS}=\frac{\text{Model\ Score}}{\text{SOTA\ Score}}.$ (5)
@@
RPS=SOTA ScoreModel $\text{RPS}=\frac{\text{SOTA\ Score}}{\text{Model\ Score}}.$ (6)
By using these normalized values, we can uniformly compare model performance across tasks where different evaluation metrics may have opposite directions (larger vs. smaller). In both cases, a value close to 1 indicates that the model’s performance is nearly on par with SOTA. This approach allows for an intuitive and consistent representation of a relative comparison of the model’s performance against the best-known performance (SOTA) in a range of 0 to 1, where 1 represents perfect or near-perfect performance on the task. Note that in our analysis, the normalized value of ASR task is the the average of two values estimated on LibriSpeech test-clean set and test-other set respectively.
TABLE V: Performance comparison of representative models across four core speech understanding tasks: Automatic Speech Recognition (ASR), Speech Translation (ST), Emotion Recognition (ER), and Human Sound Event Classification (SEC). The table reports both raw scores and each model’s relative performance compared to the task-specific SOTA (RPS). The final two columns summarize each model’s average RPS and its standard deviation across all tasks it participates in. Since SALMONN, Kimi-Audio, and OSUM are not capable of all the tasks we tested, the asterisk * in the result indicates the mean and standard deviation of SOTA score are calculated only based on the tasks they can handle.
@@ VII-B2 Performance Comparison and Key Insights
While the previous section provides a systematic analysis of current Speech LLM capabilities from the informational perspective, it primarily focuses on tasks mainly involving perception and shallow cognition. From the standpoint of cognitive depth, as illustrated in Section II-B, however, it offers limited evaluation of tasks requiring deeper reasoning or complex inference. To address this gap, several recent studies have proposed dedicated benchmarks targeting such cognitively demanding tasks and have conducted comprehensive evaluations of leading Speech LLMs. The results of these evaluations are summarized in Table VI.
As shown in Table VI, current Speech LLMs still exhibit suboptimal performance on benchmarks involving deeper levels of speech understanding. These benchmarks evaluate models on closed-ended or mixed open/closed-ended question answering tasks. While proprietary models like Gemini 2.0 Flash and GPT-4o Audio lead with accuracy scores around 60–66%, most open-source systems perform significantly worse, with many results falling below 35% [145]. This performance gap underscores the difficulty of deriving high-level semantic inferences from spoken input. Compared to perception or shallow cognitive tasks, these benchmarks require contextual reasoning, discourse-level understanding, and multimodal integration—areas where current models remain limited in capability.
@@ VIII Challenges
To further investigate the limitations observed in the performance section, we conducted a series of experiments aimed at clarifying and localizing the underlying issues. Building upon the multidimensional perspective introduced in Section II-B, we identify two major categories of challenges.
@@ VIII-A Instruction sensitivity
TABLE VII: Sensitivity of different instruction variations for ASR, S2TT, SER and GR tasks among three Speech LLMs. The test sets for the four tasks are: LibriSpeech [103] (test-clean / test-other) for ASR, LibriSpeech (test-clean) for GR, IEMOCAP [120] (session 5) for SER, and CoVoST2 [126] (en2zh test set) for S2TT.
@@
Instruction sensitivity in LLMs has been extensively explored within natural language processing (NLP), where it is widely regarded as a key indicator of model robustness [164]. Robustness refers to an LLM’s ability to maintain consistent task performance when faced with semantically equivalent but linguistically varied instructions. These variations may include character-level differences (e.g., punctuation), synonym substitution, rephrased sentence structure, or even different requested output formats.
In the audio domain, the IFEval‑Audio benchmark [165] provides a six‑dimensional perspective—Content, Capitalization, Symbol, List Structure, Length, and Format—to evaluate instruction‑following performance in end-to-end or cascaded LLMs integrated with audio modality. However, this benchmark only assesses dialogue performance, focusing on whether models preserve the semantics of chat and QA interactions, and uses LLMs to verify if responses maintain semantic consistency. While useful, this is far from sufficient for the unified speech understanding LLMs studied in this work, which require evaluation across a broader range of SLU tasks beyond conversational compliance.
Speech LLMs are expected to produce reliable outputs for SLU tasks, which require both acoustic perception and shallow semantic reasoning. While these models typically perform well on instructions encountered during training, two critical challenges emerge in real-world scenarios where the same or similar task requirements are expressed using varied instructions:
@@
These issues pose significant practical limitations. For example, in batch-processing pipelines, instruction variation can lead to operational inefficiencies and reduced reliability. To mitigate this, practitioners often resort to manual instruction selection—testing multiple phrasings and selecting the one that yields the best performance for each task. This approach, however, is both labor-intensive and inherently unstable. To pursue more consistent performance, practitioners frequently fine-tune models using parameter-efficient fine-tuning (PEFT) methods [166, 167]. While effective in stabilizing outputs, such fine-tuning typically compromises the model’s generalization capabilities, resulting in overfitting to a narrow instruction set. Additionally, PEFT introduces computational overhead and increases deployment time, thereby making the deployment pipeline more resource-intensive.
In this survey, we formally define these phenomena as key challenges for speech LLMs in speech understanding tasks. As a preliminary step, we conduct empirical evaluations to quantify their effects. To assess acoustic perception sensitivity, we employ automatic speech recognition (ASR) and gender recognition (GR) tasks. For evaluating shallow cognition sensitivity, we utilize speech-to-text translation (S2TT) and speech emotion recognition (SER).
@@
To comprehensively evaluate the instruction sensitivity of speech LLMs for SLU tasks, we first introduce instruction following rate(IFR) to assess the model’s ability to handle the challenge of instruction following failure. This metric reflects the proportion of samples in the evaluation set for which the model is able to reasonably follow the given instructions. In our experiments, we adopted concise and explicit instructions, which instruct the model to return only the task-specific answer, omitting any extraneous or conversational content. To evaluate instruction-following, we examined insertion errors in the WER calculation: if the number of insertions exceeded 2, the output was typically considered to contain extraneous content beyond the transcription—such as prefixes like “The transcription is”—or entirely omitted the target language. In such cases, the output was deemed to have failed to follow the instruction. For the S2TT task, we directly searched for specific prefixes in the model output and then conducted a secondary verification against the ground truth to determine whether the instruction was not followed. In the SER task, instruction-following was assessed by checking whether the output was strictly one of the four expected categories: “Happy”, “Sad”, “Angry”, or “Neutral”. Similarly, for the GR task, the output was required to be either “Male” or “Female”.
@@
To address the challenge of instruction-induced performance instability, we treat samples where the instruction is not followed as empty outputs and compute performance metrics accordingly. Supposing $\mathcal{D}=(\mathcal{G},\mathcal{H})$ is the paired evaluation dataset, where $\mathcal{G}=\{g_{i}\}$ denotes the set of ground-truth outputs and =$\mathcal{H}=\{h_{i}\}$ the corresponding model outputs. Let $\mathcal{F}\subseteq\mathcal{H}$ denote the subset of outputs that correctly follow the given instructions. The resulting performance under the constraint of instruction following adherence is thus referred to as
@@
In this dimension of analysis, we first construct a general instruction and then design six variant instructions along different perturbation axes. These variants are crafted to systematically examine the model’s sensitivity to changes in instruction phrasing, including alterations in punctuation, semantic complexity, case sensitivity and format requirements. The details of instructions variants are demonstrated in Appendix XI-A. The goal is to evaluate the model’s robustness and consistency in instruction following under diverse transformations with consistent SLU task requirements. Three open-source models, which are Qwen2-Audio-7B-Instruct, Qwen2.5-Omni-7B and SALMONN-14B, are evaluated. The instruction sensitivity of a model is positively correlated with the standard deviation of performance instability metric set containing $\textbf{Metric}_{\text{IF}}$ on different axes, indicating that greater variability in performance across different instructions reflects higher sensitivity to instruction phrasing:
@@
As demonstrated in Table VII, we utilize the standard errors of $\{\textbf{Metric}_{\text{IF}}\}$ to represent instruction sensitivity. To varying degrees, it is observed across all evaluated models, posing a potential challenge to the models’ fundamental perception and cognition capabilities. The results indicate that Qwen2.5-Omni-7B exhibits relatively high instruction sensitivity, with its Instruction-Following Rate (IFR) showing sharp declines in certain tasks. This may be partially attributed to its Omni architecture. In contrast, SALMONN-13B performs poorly across most tasks—except for ASR—occasionally falling below the performance of a random guesser with 50% accuracy in the GR task, likely due to consistently outputting the same label. Nevertheless, it demonstrates exceptionally low instruction sensitivity, maintaining an instruction-following rate close to 100%. We hypothesize that a model’s degree of instruction sensitivity is closely related to the types of instructions it encountered during fine-tuning and the diversity of task-specific data it is trained on.
@@ VIII-B Degradation on semantic understanding ability
TABLE VIII: Challenge B (section VIII-B): comparison of Speech LLMs and base LLMs on ST (Covost2) and SLU (MMSU). MMSU dataset does not contain GT text transcription of the speech, so the corresponding rows are emitted.
@@
Qwen2.5-7B-Instruct [92] GT Text 43.97 - - - - -
@@
Table VIII shows the results. And the detailed prompts are listed in Section XI-B. In the ST task, from the perspective of Qwen-Omni, we observe that while its overall translation performance improves compared to the Instruct variant due to better task fine-tuning, the results across the three input conditions (Speech, ASR Text, and GT Text) indicate that integrating the speech modality leads to less satisfactory end-to-end transcription results. For the SALMONN experiments, given that Vicuna lacks sufficient instruction tuning, its translation performance is noticeably inferior to the SALMONN model. Moreover, within the SALMONN series, the integration of the speech modality demonstrates relatively strong overall performance, demonstrating the effectiveness of the Speech LLM paradigm on shallow cognitive tasks.
@@
In general, these results suggest that although Speech LLMs can align speech inputs with textual outputs through fine-tuning, their original textual reasoning capacity may be weakened, likely due to modality alignment constraints and shared model capacity.
Beyond the two challenges elaborated above, Speech LLMs face a range of additional issues. These include difficulties in reliably capturing acoustic information and limited reasoning capabilities based on this, and challenges with temporal alignment of speech segments—particularly in defining independent time boundaries through instructions, an aspect that has received little attention due to the inherent coupling of alignment with other tasks. Furthermore, managing speaker turns and maintaining correct sequential order in multi-party conversations remain important areas warranting further investigation.

### v6

@@ VII-A2 Inaccurate Evaluation
In summary, our dual-track evaluation framework accommodates structured, unstructured, and weakly-structured tasks, fostering comprehensive and principled analysis of Speech LLMs across diverse speech understanding scenarios. To further ground this framework in practice, the following sections present a detailed analysis of representative tasks from each category—namely, ASR as a canonical structured task, ST as a prototypical weakly-structured task, and multitask capabilities as a hallmark of unstructured understanding.
@@ VII-B1 Analysis Setup
$\text{RPS}=\frac{\text{Model\ Score}}{\text{SOTA\ Score}}.$ (5)
@@
$\text{RPS}=\frac{\text{SOTA\ Score}}{\text{Model\ Score}}.$ (6)
By using these normalized values, we can uniformly compare model performance across tasks where different evaluation metrics may have opposite directions (larger vs. smaller). In both cases, a value close to 1 indicates that the model’s performance is nearly on par with SOTA. This approach allows for an intuitive and consistent representation of a relative comparison of the model’s performance against the best-known performance (SOTA) in a range of 0 to 1, where 1 represents perfect or near-perfect performance on the task. Note that in our analysis, the normalized value of ASR task is the average of two values estimated on LibriSpeech test-clean set and test-other set, respectively.
TABLE V: Performance comparison of representative models across four core speech understanding tasks: Automatic Speech Recognition (ASR), Speech Translation (ST), Emotion Recognition (ER), and Human Sound Event Classification (SEC). The table reports both raw scores and each model’s relative performance compared to the task-specific SOTA (RPS). The final two columns summarize each model’s average RPS and its standard deviation across all tasks it participates in. Since SALMONN, Kimi-Audio, and OSUM are not capable of all the tasks we tested, the asterisk * in the result indicates that the mean and standard deviation of SOTA score are calculated only based on the tasks they can handle.
@@ VII-B2 Performance Comparison and Key Insights
While the previous section provides a systematic analysis of current Speech LLM capabilities from the informational perspective, it primarily focuses on tasks mainly involving perception and shallow cognition. From the standpoint of cognitive depth, as illustrated in Section II-B, however, it offers limited evaluation of tasks requiring deeper reasoning or complex inference. To address this gap, several recent studies have proposed dedicated benchmarks targeting such cognitively demanding tasks and have conducted comprehensive evaluations of leading Speech LLMs. The results of these evaluations are summarized in Table VI. These benchmarks evaluate models on closed-ended or mixed open/closed-ended question answering tasks.
In addition to tasks that primarily capture acoustic content or straightforward semantic inference, these benchmarks probe cognitively richer abilities, such as emotion-state summarization from speech and multi-speaker role mapping. For example, in the MMAU benchmark, the Speech Emotion State Summarisation and Multi-Speaker Role Mapping tracks evaluate how well models integrate prosody, timbre, and speaker-dependent cues with textual reasoning. These tasks highlight that Speech LLMs must jointly model fine-grained acoustic signals (e.g., emotional valence, speaker traits, turn-taking structure) and textual semantics to achieve genuine speech cognition.
As shown in Table VI, current Speech LLMs still exhibit suboptimal performance on benchmarks involving deeper levels of speech understanding. While proprietary models like Gemini 2.0 Flash and GPT-4o Audio lead with accuracy scores around 60–66%, most open-source systems perform significantly worse, with many results falling below 35% [157]. This performance gap underscores the difficulty of deriving high-level semantic inferences from spoken input. Compared to perception or shallow cognitive tasks, these benchmarks require contextual reasoning, discourse-level understanding, and multimodal integration—areas where current models remain limited in capability.
@@
VII-C Model Performance in Real-World Scenarios
While the previous subsections evaluate Speech LLMs on standard benchmarks, real-world deployment introduces additional considerations beyond those captured in curated datasets. In practice, user speech often contains background noise, interruptions, informal phrasing, and diverse recording conditions. As a result, performance in in-the-wild scenarios may deviate from results obtained on clean academic benchmarks. For example, several recent Speech LLMs report stable performance on large-scale real-world corpora such as WenetSpeech [175] and GigaSpeech [123], suggesting that their robust acoustic front-ends and large-scale pretraining help maintain recognition quality in unconstrained settings. However, the lack of a unified and widely accepted real-world evaluation benchmark makes it difficult to directly compare models across practical application contexts.
In addition, real-world systems often need to satisfy specific deployment constraints. One such challenge arises in low-resource and multilingual scenarios, where available speech data is sparse or covers diverse accents and dialects. Current Speech LLMs often rely on high-resource pretraining and thus do not consistently outperform traditional ASR pipelines or multilingual speech models in these settings, as shown in Table VII [82, 176, 64]. Another challenge is real-time interaction. The large parameter scales of many Speech LLMs lead to high inference latency, posing difficulties for latency-sensitive applications such as interactive voice assistants, customer service agents, and dialogue-driven human–robot interaction. Although quantization, model distillation, and streaming decoding architectures have been explored to reduce latency, real-time, fully multimodal LLM-based interaction remains an active research problem.
Overall, while Speech LLMs demonstrate increasingly competitive robustness in both controlled evaluation settings and real-world corpora, their performance and practicality in truly diverse, noise-rich, multilingual, and low-latency scenarios remain limited. Addressing these challenges is essential for advancing Speech LLMs toward reliable, real-world conversational intelligence.
TABLE VII: Comparison of multilingual ASR performance on the FLEURS dataset (average WER across 19 languages).
Category Model WER $\downarrow$
E2E specialized model Whisper Large-v2 8.86
Whisper Large-v3 6.52
Open-source SLLM Qwen2.5-Omni 14.04
Qwen3-Omni-30B-A3B-Instruct 5.33
Qwen3-Omni-Flash-Instruct 5.31
Proprietary SLLM GPT-4o-Transcribe 4.48
Gemini-2.5-Pro 5.55
  • Languages included: Arabic, Cantonese, Chinese, Dutch, English, French, German, Indonesian, Italian, Japanese, Korean, Malay, Portuguese, Russian, Spanish, Thai, Turkish, Urdu, and Vietnamese.
@@ VIII Challenges
To further investigate the limitations observed in the performance section, we conducted a series of experiments aimed at clarifying and localizing the underlying issues. Building on the multidimensional perspective introduced in Section II-B, we identify two main categories of challenges.
@@ VIII-A Instruction sensitivity
TABLE VIII: Sensitivity of different instruction variations for ASR, S2TT, SER and GR tasks among three Speech LLMs. The test sets for the four tasks are: LibriSpeech [115] (test-clean / test-other) for ASR, LibriSpeech (test-clean) for GR, IEMOCAP [132] (session 5) for SER, and CoVoST2 [138] (en2zh test set) for S2TT.
@@
Instruction sensitivity in LLMs has been extensively explored within natural language processing (NLP), where it is widely regarded as a key indicator of model robustness [177, 178]. Robustness refers to the ability of LLMs to maintain consistent task performance when faced with semantically equivalent but linguistically varied instructions. These variations may include character-level differences (e.g., punctuation), synonym substitution, rephrased sentence structure, or even different requested output formats.
In the audio domain, the IFEval-Audio benchmark [179] provides a six-dimensional perspective on instruction-following performance in end-to-end or cascaded LLMs that integrate the audio modality. It evaluates models along six dimensions: Content, Capitalization, Symbol, List Structure, Length, and Format. However, this benchmark only assesses dialog performance, focusing on whether models preserve the semantics of chat and QA interactions, and uses LLMs to verify if responses maintain semantic consistency. Although useful, this is far from sufficient for the unified speech understanding LLMs studied in this work, which require evaluation across a broader range of SLU tasks beyond conversational compliance.
Speech LLMs are expected to produce reliable outputs for SLU tasks, which require both acoustic perception and shallow semantic reasoning. Although these models typically perform well on instructions encountered during training, two critical challenges emerge in real-world scenarios where the same or similar task requirements are expressed using varied instructions:
@@
These issues pose significant practical limitations. For example, in batch-processing pipelines, variation in instruction can lead to operational inefficiencies and reduced reliability. To mitigate this, practitioners often resort to manual instruction selection—testing multiple phrasings and selecting the one that yields the best performance for each task. However, this approach is labor-intensive and inherently unstable. To achieve more consistent performance, practitioners frequently fine-tune models using parameter-efficient fine-tuning methods (PEFT)  [180, 181]. Although effective in stabilizing outputs, this fine-tuning typically compromises the model’s generalization capabilities, resulting in overfitting to a narrow instruction set. Additionally, PEFT introduces computational overhead and increases deployment time, making the deployment pipeline more resource-intensive.
In this survey, we formally define these phenomena as key challenges for speech LLMs in speech understanding tasks. As a preliminary step, we conduct empirical evaluations to quantify their effects. To assess acoustic perception sensitivity, we employ automatic speech recognition (ASR) and gender recognition (GR) tasks. To evaluate shallow cognition sensitivity, we utilize speech-to-text translation (S2TT) and speech emotion recognition (SER).
@@
To comprehensively evaluate the sensitivity of speech LLMs for SLU tasks, we first introduce instruction following rate(IFR) to assess the model’s ability to handle the challenge of instruction following failure. This metric reflects the proportion of samples in the evaluation set for which the model is able to reasonably follow the given instructions. In our experiments, we adopted concise and explicit instructions which instruct the model to return only the task-specific answer, omitting any extraneous or conversational content. To assess instruction-following ability, we used insertion errors in the WER calculation as an indicator. Outputs with more than 2 insertions usually contained extraneous content beyond the transcription, for example prefixes like “The transcription is”, or even omitted the target language altogether. In such cases, the output was deemed to have failed to follow the instruction. For the S2TT task, we directly searched for specific prefixes in the model output and then conducted a secondary verification against the ground truth to determine whether the instruction was not followed. In the SER task, the follow up of the instruction was assessed by checking whether the output was strictly one of the four expected categories: “Happy”, “Sad”, “Angry”, or “Neutral”. Similarly, for the GR task, the output was required to be either “Male” or “Female”.
@@
To address the challenge of instruction-induced performance instability, we treat samples where the instruction is not followed as empty outputs and compute performance metrics accordingly. Suppose $\mathcal{D}=(\mathcal{G},\mathcal{H})$ is the paired evaluation dataset, where $\mathcal{G}=\{g_{i}\}$ denotes the set of ground-truth outputs and $\mathcal{H}=\{h_{i}\}$ the corresponding model outputs. Let $\mathcal{F}\subseteq\mathcal{H}$ denote the subset of outputs that correctly follow the given instructions. The resulting performance under the constraint of instruction following adherence is thus referred to as
@@
In this dimension of analysis, we first construct a general instruction and then design six variant instructions along different perturbation axes. These variants are designed to systematically examine the model’s sensitivity to changes in instruction phrasing, including alterations in punctuation, semantic complexity, case sensitivity, and format requirements. The details of the instructions variants are demonstrated in Appendix XI-A. The goal is to evaluate the robustness and consistency of the instruction following in diverse transformations with consistent SLU task requirements. Three open-source models are evaluated, which are Qwen2-Audio-7B-Instruct, Qwen2.5-Omni-7B and SALMONN-14B. The instruction sensitivity of a model is positively correlated with the standard deviation of the performance instability metric set containing $\textbf{Metric}_{\text{IF}}$ on different axes, indicating that greater variability in performance across different instructions reflects higher sensitivity to instruction phrasing:
@@
As demonstrated in Table VIII, we utilize the standard errors of $\{\textbf{Metric}_{\text{IF}}\}$ to represent instruction sensitivity. To varying degrees, it is observed across all evaluated models, which poses a potential challenge to the models’ fundamental perception and cognition capabilities. The results indicate that Qwen2.5-Omni-7B exhibits relatively high instruction sensitivity, with its Instruction-Following Rate (IFR) showing sharp declines in certain tasks. This may be partially attributed to its Omni architecture. In contrast, SALMONN-13B performs poorly on most tasks, with the exception of ASR. On the GR task, its accuracy occasionally falls below that of a random guesser with 50% accuracy, likely because it consistently outputs the same label. Nevertheless, it demonstrates exceptionally low instruction sensitivity, maintaining an instruction-following rate close to 100%. We hypothesize that a model’s degree of instruction sensitivity is closely related to the types of instructions it encountered during fine-tuning and the diversity of task-specific data it is trained on.
@@ VIII-B Degradation on semantic understanding ability
TABLE IX: Challenge B (section VIII-B): comparison of Speech LLMs and base LLMs on ST (Covost2) and SLU (MMSU). MMSU dataset does not contain GT text transcription of the speech, so the corresponding rows are emitted.
@@
Qwen2.5-7B-Instruct [104]
GT Text 43.97 - - - - -
@@
Table IX shows the results. And the detailed prompts are listed in Section XI-B. In the ST task, from the perspective of Qwen-Omni, we observe that while its overall translation performance improves compared to the Instruct variant due to better task fine-tuning, the results across the three input conditions (Speech, ASR Text, and GT Text) indicate that integrating the speech modality leads to less satisfactory end-to-end transcription results. For the SALMONN experiments, given that Vicuna lacks sufficient instruction tuning, its translation performance is noticeably inferior to the SALMONN model. Moreover, within the SALMONN series, the integration of the speech modality demonstrates relatively strong overall performance, demonstrating the effectiveness of the Speech LLM paradigm on shallow cognitive tasks.
@@
In general, these results suggest that although Speech LLMs can align speech inputs with textual outputs through fine-tuning, their original textual reasoning capacity may be weakened, likely due to modality alignment constraints [64] and shared model capacity.
Beyond the two challenges discussed above, Speech LLMs encounter several additional limitations. First, they often struggle to reliably capture and reason over acoustic information, leading to restricted performance in tasks that require fine-grained auditory understanding. Another challenge lies in the temporal alignment of speech segments, in particular the definition of independent time boundaries using instruction-based mechanisms. This aspect has been relatively underexplored, as it is tightly coupled with downstream tasks. Finally, managing speaker turns and preserving correct sequential order in multi-party conversations remain open problems that warrant further investigation.